# ALPINE: Unveiling the Planning Capability of Autoregressive Learning in Language Models

### v2

@@ 2.2 Path Planning Dataset
The dataset is designed to test GPT’s path planning capability on simple graphs. We consider a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes, and $\mathcal{E}$ is the set of directed edges, i.e., for any $u,v\in\mathcal{V}$, $(u,v)\in\mathcal{E}$ means that there is an edge from node $u$ to node $v$ in $\mathcal{G}$. A pair of source node $s$ and target node $t$ is considered as a valid pair if $\mathcal{G}$ contains least one path from $s$ to $t$. We allocate a portion of valid $(s,t)$ pairs to the training dataset and assign the remaining pairs to the test dataset.
@@
3 Path-Finding Transformers: Expressive Power and Learning Capacity
@@
2:Output: A list contains a path from $s$ to $t$
3:Set path node $P=[s]$ and set current node $u=s$
4:while $i\neq t$ do
5:     Obtain $S=\{k|\bm{A}_{(i,k)}=1\text{ and }\bm{R}_{(t,k)}=1\}$
6:     Randomly sample next node $k$ from $S$
7:     Add $k$ to the results $P=[P;k]$
8:     Set $i=k$
9:end while
In this section, we present an overview of our main results. Firstly, in Theorem 2 below, we establish the mathematical existence of a Transformer model capable of effective path finding in any given network. Next, in Theorem 3— the proof of which will be the main focus of Section 4—we characterize the learning capacity and limitations of path-finding transformer models in a data-driven gradient descent framework. The empirical evaluation that supports these theoretical analyses will be discussed in Section 5.
In our path-finding task, the essential step for completing a path is to predict the next node based on the current information. It is apparent that to predict the subsequent node on the path, only the information related to the current node and the target node is necessary. Algorithm 1 introduces a handcrafted algorithm that utilizes both the adjacency matrix and the reachability matrix of the graph. The true adjacency matrix follows the standard definition in graph theory, i.e.,
$\bm{A}^{\text{\rm true}}_{(i,k)}=\begin{cases}1,&\text{if }(i,k)\in\mathcal{E},\\ 0,&\text{otherwise.}\end{cases}$
The true reachability matrix is defined as:
$\bm{R}^{\text{\rm true}}_{(t,k)}=\begin{cases}1,&\text{if }k\text{ can reach }t\text{ in }\mathcal{G},\\ 0,&\text{otherwise.}\end{cases}$
@@
Assuming that $t$ is reachable by $s$, then Algorithm 1 is guaranteed to output a valid path with input $\bm{A}=\bm{A}^{\text{\rm true}}$ and $\bm{R}=\bm{R}^{\text{\rm true}}$.
To Illustrate the expressive capacities of the Transformer model in path finding, we first show that we can manually construct a Transformer model to perform the path planning task by simulating the idealized Algorithm 1. In the manual construction, the task for the model is to find a path from a start node $s$ to a target node $t$ with the format “$s\ t\ s\ u_{1}\ u_{2}\ \ldots\ u_{p}\ t$”. Consider every time that the Transformer takes “$s\ t\ s\ u_{1}\ u_{2}\ \ldots\ u_{k}$” as the input and outputs $u_{k+1}$ for $k=0,1,\ldots,p$ (assuming $s=u_{0}$ and $t=u_{p+1}$): if $u_{k+1}$ is an out-neighbor of $u_{k}$ and can reach $t$ in $\mathcal{G}$, then we say that the Transformer outputs a correct response.
Theorem 2 (Expressive Power of the Transformer Model in Path Finding).

### v3

@@ 2.2 Path-Planning Dataset: Syntax and Data Sources
The dataset is designed to evaluate GPT’s path-planning capability on simple graphs. It is generated from a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes, and $\mathcal{E}$ is the set of edges. For any two nodes $u,v\in\mathcal{V}$, $(u,v)\in\mathcal{E}$ means that there is a directed edge from $u$ to $v$ in $\mathcal{G}$. A pair of source node $s$ and target node $t$ is considered a valid pair if $\mathcal{G}$ contains at least one path from $s$ to $t$.
@@
3 Expressiveness and Learning Dynamics of Transformer Models
3.1 Expressiveness
@@
2:Set path $P=[s\ t\ s]$ and set current node $i=s$
3:while $i\neq t$ do
4:     Obtain $S=\{k|\bm{A}_{(i,k)}=1\text{ and }\bm{R}_{(t,k)}=1\}$
5:     Randomly sample next node $k$ from $S$
6:     Append $k$ to path $P$, and set $i=k$
7:end while
8:output path $P$
In our path-finding task, the essential step for completing a path is to predict the next node based on the current information. It is evident that to predict the subsequent node on the path, only information related to the current node and the target node is necessary. Algorithm 1 introduces a idealized method that utilizes both the adjacency and reachability matrices of the graph. The true adjacency matrix $\bm{A}^{\text{\rm true}}$ and the true reachability matrix $\bm{R}^{\text{\rm true}}$ are defined as:
$\bm{A}^{\text{\rm true}}_{(i,k)}=\begin{cases}1,&\text{if }(i,k)\in\mathcal{E},\\ 0,&\text{otherwise.}\end{cases}\qquad\bm{R}^{\text{\rm true}}_{(t,k)}=\begin{cases}1,&\text{if }k\text{ can reach }t\text{ in }\mathcal{G},\\ 0,&\text{otherwise.}\end{cases}$
@@
Assuming that $t$ is reachable from $s$, then Algorithm 1 is guaranteed to output a valid path with input $\bm{A}=\bm{A}^{\text{\rm true}}$ and $\bm{R}=\bm{R}^{\text{\rm true}}$.
To illustrate the expressive capacities of the Transformer model, we first demonstrate how to manually construct a Transformer that can perform the path-finding task by simulating Algorithm 1.
Theorem 2.